# CountZES: Counting via Zero-Shot Exemplar Selection

### v1

@@ 3.2 Density-Guided Exemplar (DGE)
Peak-to-Point (P2P) prompting: Given the input image $I$ and exemplar $\mathbf{b}_{\text{DAE}}$, the pretrained counter produces a density map $D\in\mathbb{R}^{H\times W}$, where each pixel value estimates the expected object count (Fig. 3). The total count is $C=\sum_{x,y}D(x,y)$, and local maxima in $D$ correspond to potential object centers. Assuming $D$ follows an approximately Gaussian distribution, most pixel responses correspond to background while object peaks appear as statistical outliers. Thus, an adaptive threshold is defined as $T_{D}=\mu_{D}+2\sigma_{D}$, where $\mu_{D}$ and $\sigma_{D}$ are the mean and standard deviation of $D$. Pixels satisfying $D(x,y)\geq T_{D}$ are retained as high-confidence candidates. Distinct peaks are then identified via local peak detection, which retains only the strongest responses within a fixed neighborhood, yielding a set of spatially separated candidate points:
@@
To ensure semantic consistency with GroundingDINO detections, we apply box gating, retaining only peaks $p_{i}$ that fall within at least one detection box $\mathbf{b}j^{\text{DINO}}$. The filtered prompt set is thus $\mathcal{P}^{\star}=\bigl\{\,p_{i}\in\mathcal{P}\mid p_{i}\in\Omega_{\text{DINO}}\,\bigr\}$, where $\Omega_{\text{DINO}}$ denotes the union of all detection boxes. Each surviving point $p_{i}\in\mathcal{P}^{\star}$ is used as a positive prompt for SAM, generating a binary mask $M_{i}=\mathcal{P}(I;\,p_{i})$ and its corresponding bounding box $\mathbf{b}_{i}$. These boxes then serve as candidate exemplars for subsequent selection.
@@
where $\Omega(\mathbf{b}_{i})$ denotes the set of pixel coordinates inside the box. This operation is conceptually inspired by the RoI pooling mechanism in Fast R-CNN [girshick2015fast]. Since regression noise and density diffusion may yield fractional counts above one, boxes with predicted counts satisfying $1<\hat{c}(\mathbf{b}i)<2$ are retained as single-instance candidates. The resulting subset $\mathcal{B}_{\text{single}}$ represents single-instance proposals for subsequent exemplar refinement.
Figure 4: Overview of GGES module in the DGE stage. Per-box counts from single-instance candidates are used to estimate a pseudo-GT count via non-parametric density estimation. A SAM-based similarity map, conditioned on the DAE exemplar, captures semantic correspondence, and a composite score combining count proximity to the pseudo-GT and similarity-map entropy identifies the final density-guided exemplar.
Pseudo-GT guided exemplar selection (GGES): To identify the most reliable exemplar within $\mathcal{B}_{\text{single}}$, we estimate a pseudo-GT count that reflects the consensus among all single-instance candidates. Rather than relying on discrete per-box predictions $\hat{c}_{i}=\hat{c}(\mathbf{b}_{i})$}, we model their empirical distribution using a non-parametric probability density function (PDF). This continuous PDF reveals the dominant count region, consolidating multiple uncertain predictions into a stable pseudo-GT reference (Fig. 4). By leveraging the overall distribution instead of individual estimates, this approach mitigates outlier effects and guides exemplar selection toward the statistically most consistent estimate.
Each candidate box $\mathbf{b}_{i}$ is then evaluated for its closeness to the pseudo-GT count, defined as $C_{\text{close}}(\mathbf{b}_{i})$. Higher $C_{\text{close}}(\mathbf{b}_{i})$ indicate higher representativeness as a density-guided single-instance exemplar. However, count consistency alone does not ensure semantic purity. A region may agree with pseudo-GT yet still include clutter or background. To assess the semantic compactness of each candidate region, we compute an exemplar-conditioned SAM similarity map. Unlike the DAE stage, where a text-conditioned CLIP similarity map was used due to the absence of a concrete exemplar, the DGE stage benefits from having the previously obtained $\mathbf{b}_{\text{DAE}}$ exemplar. This enables a more discriminative and instance-specific similarity computation. Specifically, $\mathbf{b}_{\text{DAE}}$ is passed through SAM’s encoder to obtain its feature embedding, which is then compared, via cosine similarity, with the spatial embeddings of the entire image. The resulting similarity map provides a more discriminative and instance-specific representation, highlighting regions that share strong visual and structural correspondence with the exemplar (Fig. 4).
@@
where $\alpha=0.5$ balances count consistency and semantic compactness. The box with the highest score is selected as DGE, $\mathbf{b}_{\text{DGE}}=\arg\max_{\mathbf{b}_{i}\in\mathcal{B}_{\text{single}}}J_{\text{DGE}}(\mathbf{b}_{i})$. DGE thus complements the text-conditioned DAE by introducing a count-driven self-supervised perspective. By grounding exemplar selection in the predicted density rather than textual cues, DGE enhances coverage and robustness, particularly in cluttered or repetitive scenes where text-conditioned grounding alone is insufficient.
@@ 3.3 Feature-Consensus Exemplar (FCE)
While DGE enforces statistical consistency via pseudo-GT estimation, it does not explicitly capture visual coherence among candidate instances. The FCE stage refines exemplar selection at the feature level, ensuring that the final exemplar is visually representative of the dominant object appearance within the scene. Unlike DAE, which begins from text-conditioned grounding, FCE builds upon the exemplar from the DGE stage, $\mathbf{b}_{\text{DGE}}$. Using $\mathbf{b}_{\text{DGE}}$ as the anchor, FCE regenerates a density map, and performs P2P prompting followed by single-instance filtering, mirroring the steps of DGE up to the formation of $\mathcal{B}_{\text{single}}$. However, instead of pseudo-GT guided selection, FCE introduces a Feature-based Representative Exemplar Selection (FRES).
Feature-based Rep Exemplar Selection (FRES): Given the set of single-instance boxes $\mathcal{B}_{\text{single}}$, our goal is to identify the exemplar that most effectively captures the dominant visual appearance of the object class (Fig. 5). The image $I$ is first passed through SAM’s encoder to obtain a dense feature map $\Phi\in\mathbb{R}^{256\times 64\times 64}$. Since projecting small boxes onto this low-resolution map may cause spatial detail loss, $\Phi$ is upsampled to $\Phi^{\text{up}}\in\mathbb{R}^{256\times 256\times 256}$ using window attention–based upsampling as in [wimmer2025anyup]. Each candidate box $\mathbf{b}_{i}$ is then projected onto $\Phi^{\text{up}}$, and the corresponding region $\Phi^{\text{up}}(\Omega_{f}(\mathbf{b}_{i}))$ defines its local feature subset. Finally, each box descriptor $\mathbf{v}_{i}$ is computed via spatial average pooling:
@@
The vector $\mathbf{v}_{i}$ thus represents the semantic appearance of region $\mathbf{b}_{i}$ in the SAM feature space. Since the magnitude of pooled features may vary with object size, texture, or illumination, direct comparison between descriptors can be biased toward high-energy regions. To mitigate this effect, we apply $\ell_{2}$ normalization so that all descriptors lie on a unit hypersphere, ensuring that their pairwise dot products correspond to cosine similarity.
While each descriptor $\mathbf{v}_{i}$ captures the visual characteristics of a single-instance region, these descriptors may still vary due to pose, scale, illumination, or minor background interference. To isolate the most representative exemplar, we cluster the descriptors in SAM’s feature space to identify the dominant appearance pattern. Specifically, we cluster $\mathbf{v}_{i}$} into two groups using cosine similarity, forming a majority and a minority cluster. This partition effectively separates the visually consistent, frequently occurring appearances from peripheral or noisy ones. The majority cluster $\mathcal{C}_{\text{maj}}$ captures the dominant object appearance, and the exemplar whose descriptor is closest to its centroid is selected as the representative exemplar, $\mathbf{b}_{\text{FCE}}=\arg\max_{\mathbf{b}_{i}\in\mathcal{C}_{\text{maj}}}s(\mathbf{v}_{i},\mu_{\text{maj}})$. Where $s(\cdot,\cdot)$ denotes cosine similarity and $\mu_{\text{maj}}$ is the normalized mean feature of the majority cluster. The selected exemplar $\mathbf{b}_{\text{FCE}}$ thus reflects the shared visual characteristics across all single-instance proposals. By grounding exemplar selection in feature-level consensus rather than count statistics, FCE complements DGE by emphasizing visual coherence.
Ultimately, the best exemplars obtained from DAE, DGE, and FCE collectively form the final exemplar set for counting. Selecting one exemplar per stage, rather than multiple exemplars from a single stage, introduces cross-perspective diversity (see ablations in Sec. 4). This diversity improves counting performance by enabling the counter to perceive object variations more comprehensively and reducing bias toward any single representation modality.
@@ 4 Experiments
Datasets and Evaluation Metrics. We evaluate our CountZES framework on diverse benchmarks including FSC-147 [shi2022represent], CARPK [hsieh2017drone], PerSense-D [siddiqui2024persense], and two medical cell-counting datasets namely, VGG [lempitsky2010learning] and MBM [paul2017count]. FSC-147 dataset contains 6,135 paired images across 147 object classes. CARPK dataset provides a bird’s-eye view of 89,777 cars across 1,448 parking-lot images, serving as a standard benchmark for assessing cross-domain generalization. The recently introduced PerSense-D dataset serves as a dense counting benchmark, comprising 717 images spanning 28 object categories, with an average of 53 target instances per image. The VGG dataset comprises 200 fluorescence microscopy images with simulated bacterial cells captured at varying focal depths. The MBM dataset includes 44 microscopic images containing 126 ± 33 nuclei per image, posing a challenging counting scenario due to high appearance variability. Following prior class-agnostic object counting works [qian2025t2icount, xu2023zero, zhu2024zero], we employ mean absolute error (MAE) to measure counting accuracy and root mean square error (RMSE) to assess model robustness.
@@
Implementation Details. CountZES employs CLIP [radford2021learning] for text–image alignment and GroundingDINO [liu2024grounding] for initial box proposals, with the detection threshold set to 0.15. SAM [kirillov2023segment] is used for mask generation leveraging point prompts across the SSES, GGES, and FRES modules. To demonstrate the model-agnostic nature of CountZES, we separately utilize CounTR [liu2022countr] and DSALVANet [he2024few] as pretrained off-the-shelf few-shot object counters. When the counter is pretrained on the same dataset as the one being evaluated, we denote this configuration as in-domain. Under the training-free setting, CountZES employs counter pretrained on a dataset different from the one being tested. Specifically, when evaluating on FSC-147, we use a counter pretrained on CARPK, whereas for all remaining datasets, the counter pretrained on FSC-147 is employed.
Quantitative Result on FSC-147. We evaluate CountZES under two distinct configurations as shown in Table 1. In the in-domain setting, DSALVANet pretrained on FSC-147 dataset is used as the counting backbone. Under this setup, CountZES achieves the best RMSE among SOTA while maintaining a competitive MAE. Compared to the leading exemplar selection–based method, VA-Count [zhu2024zero], CountZES lowers MAE by 11.8% and RMSE by 29.30%, indicating that the exemplars identified by our method are of higher quality. In the training-free setting, CountZES achieves SOTA performance with lowest MAE and RMSE.
@@
Table 1: Comparison of CountZES with SOTA ZOC methods on FSC-147 dataset. Best results are in bold and second-best are underlined. indicates ZOC methods based on exemplar discovery.
[…]
ZSOC [xu2023zero] CVPR’23 Yes 22.09 115.17
CLIP-Count [jiang2023clip] ACM’23 Yes 17.78 106.62
CounTX [amini2023open] BMVC’23 Yes 15.88 106.29
VLCounter [kang2024vlcounter] AAAI’24 Yes 17.05 106.16
PseCo [huang2024point] CVPR’24 Yes 16.58 129.77
DAVE [pelhan2024dave] ICCV’24 Yes 14.90 103.42
VA-Count [zhu2024zero] ECCV’24 Yes 17.88 129.31
GeCo [pelhan2024novel] NeurIPS’24 Yes 13.30 108.72
T2ICount [qian2025t2icount] CVPR’25 Yes 11.76 97.86
[…]
SAM [kirillov2023segment] ICCV’23 No 42.48 137.50
Count Anything [ma2023can] arXiv’23 No 27.97 131.24
GroundingDINO [liu2024grounding] ECCV’24 No 59.23 159.28
TFOC [shi2024training] WACV’24 No 24.79 137.15
TFCAC [lin2025simple] WACV’25 No 23.59 113.60
[…]
@@
[…]
CLIP-Count [jiang2023clip] 11.96 16.61
CounTX [amini2023open] 8.13 10.87
GroundingDINO [liu2024grounding] 29.72 31.60
VA-Count [zhu2024zero] 10.63 13.20
GeCo [pelhan2024novel] 10.34 14.73
T2ICount [qian2025t2icount] 8.61 13.47
[…]
Quantitative Result on CARPK. To evaluate cross-dataset generalization, we test CountZES on CARPK using DSALVANet as a counter pretrained on FSC-147. CountZES attains the lowest MAE and RMSE (Table 2), demonstrating remarkable transferability across domains (natural vs aerial). Notably, CountZES outperforms T2ICount [qian2025t2icount] in generalization even though it is the current SOTA on FSC-147 dataset. This superior generalization can be attributed to our training-free design, which avoids dataset-specific bias typically introduced through training.
@@
[…]
GeCo [pelhan2024novel] 11.63 17.84 10.30 20.61 12.38 24.13 13.12 28.36
T2ICount [qian2025t2icount] 30.04 122.91 25.16 76.95 20.62 52.23 24.95 88.21
[…]
@@
[…]
GeCo [pelhan2024novel] 67.32 77.53 92.59 104.52
T2ICount [qian2025t2icount] 151.35 162.81 30.01 44.44
[…]
Figure 6: Qualitative comparison of CountZES with T2ICount [qian2025t2icount] and GeCo [pelhan2024novel] across diverse scenarios, including natural (FSC-147 [shi2022represent] and PerSense-D [siddiqui2024persense]), aerial (CARPK [hsieh2017drone]), and medical cell-counting (VGG [lempitsky2010learning] and MBM [paul2017count]) images.
Quantitative Result on PerSense-D. Following the same cross-dataset generalization protocol, we evaluate CountZES on the PerSense-D dataset (Table 3). In addition to overall category-wise performance, we conduct fine-grained analysis across three density-based splits; low, medium, and high, as defined in [siddiqui2024persense]. Compared to SOTA, CountZES demonstrates competitive performance under varying crowding conditions. Notably, 16 out of 28 classes ($\approx$ 57%) in PerSense-D overlap with FSC-147 but still T2ICount, which achieves SOTA performance on FSC-147, struggles on PerSense-D. This underscores the dataset-specific bias induced by trained models, further emphasizing our inherent generalization advantage.
Quantitative Results on VGG and MBM. To assess the cross-domain generalization capability of CountZES, we evaluate it on two medical-domain cell counting datasets: VGG and MBM. As illustrated in Table 4, CountZES achieves remarkable cross-domain performance, outperforming SOTA methods by a substantial margin.
@@
[…]
Qualitative Result. Fig. 6 presents a qualitative comparison between CountZES and SOTA methods on diverse images, spanning natural, aerial, and medical domains. Fig. 8 further illustrates the performance of CountZES on images from FSC-147 dataset.
Ablation Study. CountZES comprises three sequential stages in its exemplar selection pipeline. Since each stage builds upon the exemplar generated by the previous one, we perform a progressive ablation to quantify the incremental contribution of each stage toward overall counting performance. As shown in Table 5, adding successive stages consistently improves accuracy, validating their complementary roles. Finally, to examine the diversity induced by each stage in exemplar selection, we conduct another ablation under a consistent 3-exemplar setting by selecting the top three exemplars from each stage and comparing them against the standard CountZES configuration that uses one exemplar per stage (Table 5). The latter achieves superior performance (Fig. 7), indicating that cross-stage collaboration yields a richer, complementary, and more balanced exemplar set than any stage in isolation.
Figure 8: Qualitative results of CountZES on FSC-147 dataset.
Figure 10: Additional failure cases highlighting CountZES selection of incomplete boxes as exemplars. For “nail polishes,” SAM-based point prompt segmentation in CountZES captures only partial objects; for “comic books,” point prompts isolate characters or cover details instead of the full book; for “eggs,” SAM segments only the yolk region; and for “sauce bottles,” it segments only the label instead of the whole bottle. These examples show how imperfect masks can hamper downstream exemplar selection in CountZES.
Limitation. While CountZES demonstrates better generalization across domains, certain limitations stem from the use of off-the-shelf pre-trained components like object counters and SAM. In cross-domain settings, where the pretrained counter is trained on distributions different from the target domain, this can lead to noise in peak detection via P2P prompting. Also, as vanilla SAM lacks semantic awareness, a point prompt may yield only partial masks, capturing fragments of the object rather than complete instances. Although the scoring-based selection mechanism generally favors the most reliable exemplar at each stage, residual noise can still propagate when the majority of candidate exemplars are imperfect. As illustrated in Fig. 9 for the class “watches,” a point placed on the watch frequently segments only the dial while excluding the wristband, resulting in incomplete exemplar boxes. Additional examples in Fig. 10 highlight similar failures: for “nail polishes,” the mask captures only part of the bottle; for “comic books,” the prompt segments characters on the cover instead of the whole book; for “eggs,” SAM isolates only the yolk region; and for “sauce bottles,” it focuses on the label instead of the full object. Although the scoring mechanism generally prefers the best available candidates, residual noise can still propagate when most exemplars are imperfect. Nonetheless, a key strength of CountZES lies in its modular architecture. Each component, counter or segmentation backbone, can be seamlessly replaced with more advanced alternatives without altering the overall pipeline. This flexibility opens promising avenues for further enhancing cross-domain robustness and improving exemplar fidelity in future extensions.

### v2

@@ 3.2 Density-Guided Exemplar (DGE)
Peak-to-Point (P2P) prompting: Given the input image $I$ and exemplar $\mathbf{b}_{\text{DAE}}$, the density estimator produces a density map $D\in\mathbb{R}^{H\times W}$, where each pixel value estimates the expected object count (Figure. 3). The total count is $C=\sum_{x,y}D(x,y)$, and prominent local maxima in $D$ correspond to potential object centers. Rather than assuming a specific distribution for $D$, we employ the mean and standard deviation of the density map as adaptive scale statistics to identify high-confidence peaks. Specifically, an adaptive threshold is defined as $T_{D}=\mu_{D}+2\sigma_{D}$, where $\mu_{D}$ and $\sigma_{D}$ denote the mean and standard deviation of $D$, respectively. This criterion highlights salient density responses while suppressing background noise, even when the underlying distribution is skewed. Pixels satisfying $D(x,y)\geq T_{D}$ are retained as high-confidence candidates. Distinct peaks are then extracted via local peak detection within a fixed neighborhood to ensure spatial separation:
@@
To ensure semantic consistency with GroundingDINO detections, we apply box gating, retaining only peaks $p_{i}$ that fall within at least one detection box $\mathbf{b}j^{\text{DINO}}$. The filtered prompt set is thus $\mathcal{P}^{\star}=\bigl\{\,p_{i}\in\mathcal{P}\mid p_{i}\in\Omega_{\text{DINO}}\,\bigr\}$, where $\Omega_{\text{DINO}}$ denotes the union of all detection boxes. Each surviving point $p_{i}\in\mathcal{P}^{\star}$ is used as a positive prompt for SAM, generating a binary mask $M_{i}=\mathcal{P}(I;\,p_{i})$ and its corresponding bounding box $\mathbf{b}_{i}$. These boxes then serve as candidate exemplars for subsequent selection (see Appendix. A.2).
@@
where $\Omega(\mathbf{b}_{i})$ denotes the set of pixel coordinates inside the box. This operation is conceptually inspired by the RoI pooling mechanism in Fast R-CNN (Girshick, 2015). Since regression noise and density diffusion may yield fractional counts above one, boxes with predicted counts satisfying $1<\hat{c}(\mathbf{b}i)<2$ are retained as single-instance candidates. The resulting subset $\mathcal{B}_{\text{single}}$ represents single-instance proposals for subsequent exemplar refinement (see Figure. 10).
Figure 4: Overview of GGES in the DGE stage. Per-box counts from single-instance candidates are used to estimate a pseudo-GT count via density estimator. A SAM-based similarity map, conditioned on $\mathbf{b}_{\text{DAE}}$, captures semantic correspondence, and a composite score based on count proximity to the pseudo-GT and similarity-map entropy guides density-guided exemplar selection.
Pseudo-GT guided exemplar selection (GGES): To identify the most reliable exemplar within $\mathcal{B}_{\text{single}}$, we estimate a pseudo-GT count via consensus among all single-instance candidates. Rather than relying on discrete per-box predictions $\hat{c}_{i}=\hat{c}(\mathbf{b}_{i})$}, we model their empirical distribution using a non-parametric probability density function (PDF), (see Appendix. A.3). This PDF reveals the dominant count region, consolidating multiple uncertain predictions into a stable pseudo-GT reference (Figure. 4). By leveraging the overall distribution instead of individual estimates, this approach mitigates outlier effects and guides exemplar selection toward the statistically most consistent estimate.
Each candidate box $\mathbf{b}_{i}$ is then evaluated for its predicted count closeness to the pseudo-GT count, defined as $C_{\text{close}}(\mathbf{b}_{i})$. Higher $C_{\text{close}}(\mathbf{b}_{i})$ indicate higher representativeness as a density-guided single-instance exemplar. However, count consistency alone does not ensure semantic purity. A region may agree with pseudo-GT yet still include clutter or background. To assess the semantic compactness of each candidate region, we compute an exemplar-conditioned SAM similarity map. Unlike the DAE stage, where a text-conditioned CLIP similarity map was used due to the absence of a concrete exemplar, the DGE stage benefits from having the previously obtained $\mathbf{b}_{\text{DAE}}$ exemplar. This enables a more discriminative and instance-specific similarity computation. Specifically, $\mathbf{b}_{\text{DAE}}$ is passed through SAM’s encoder to obtain its feature embedding, which is then compared, via cosine similarity, with the spatial embeddings of the entire image. The resulting similarity map provides a more discriminative and instance-specific representation, highlighting regions that share strong visual and structural correspondence with the exemplar (Figure. 4).
@@
where $\alpha=0.5$ balances count consistency and semantic compactness. The box with highest score is selected as DGE, given as $\mathbf{b}_{\text{DGE}}=\arg\max_{\mathbf{b}_{i}\in\mathcal{B}_{\text{single}}}J_{\text{DGE}}(\mathbf{b}_{i})$. DGE thus complements DAE by introducing a count-driven self-supervised perspective. By grounding exemplar selection in the predicted density rather than textual cues, DGE enhances coverage and robustness, particularly in cluttered scenes where text-conditioned grounding alone is insufficient.
@@ 3.3 Feature-Consensus Exemplar (FCE)
While DGE enforces statistical consistency via pseudo-GT estimation, it does not explicitly capture visual coherence among candidate instances. The FCE stage refines exemplar selection at the feature level, ensuring visual representativeness of the dominant object appearance within the scene. Unlike DAE, which begins from text-conditioned grounding, FCE builds upon the exemplar from the DGE stage, $\mathbf{b}_{\text{DGE}}$. Using $\mathbf{b}_{\text{DGE}}$ as the anchor, FCE regenerates a density map, and performs P2P prompting followed by single-instance filtering, mirroring the steps of DGE up to the formation of $\mathcal{B}_{\text{single}}$. However, instead of pseudo-GT guided selection, FCE introduces FRES module discussed next.
Feature-based Rep Exemplar Selection (FRES): Given $\mathcal{B}_{\text{single}}$, our goal is to identify the exemplar that most effectively captures the dominant visual appearance of the object class (Fig. 5). The image $I$ is first passed through SAM’s encoder to obtain a dense feature map $\Phi\in\mathbb{R}^{256\times 64\times 64}$. Since projecting small boxes onto this low-resolution map may cause spatial detail loss, $\Phi$ is upsampled to $\Phi^{\text{up}}\in\mathbb{R}^{256\times 256\times 256}$ using window attention–based upsampling as in (Wimmer et al., 2025). Each candidate box $\mathbf{b}_{i}$ is then projected onto $\Phi^{\text{up}}$, and the corresponding region $\Phi^{\text{up}}(\Omega_{f}(\mathbf{b}_{i}))$ defines its local feature subset. Finally, each box descriptor $\mathbf{v}_{i}$ is computed via spatial average pooling:
@@
The vector $\mathbf{v}_{i}$ thus represents the semantic appearance of region $\mathbf{b}_{i}$ in the SAM feature space. Since the magnitude of pooled features may vary with object size, texture, or illumination, direct comparison between descriptors can be biased toward high-energy regions. To mitigate this effect, we apply $\ell_{2}$-normalization so that all descriptors lie on a unit hypersphere, ensuring that their pairwise dot products correspond to cosine similarity.
While each descriptor $\mathbf{v}_{i}$ captures the visual characteristics of a single-instance region, these descriptors may still vary due to pose, scale, illumination, or minor background interference. To isolate the most representative exemplar, we cluster the descriptors in feature space to identify the dominant appearance pattern. Specifically, we cluster $\mathbf{v}_{i}$} into two groups using cosine similarity, forming a majority and a minority cluster. This partition effectively separates the visually consistent, frequently occurring appearances from peripheral or noisy ones. The majority cluster $\mathcal{C}_{\text{maj}}$ captures the dominant object appearance, and the exemplar whose descriptor is closest to its centroid is selected as the representative exemplar, $\mathbf{b}_{\text{FCE}}=\arg\max_{\mathbf{b}_{i}\in\mathcal{C}_{\text{maj}}}s(\mathbf{v}_{i},\mu_{\text{maj}})$. Where $s(\cdot,\cdot)$ denotes cosine similarity and $\mu_{\text{maj}}$ is the mean feature descriptor of the majority cluster. The selected exemplar $\mathbf{b}_{\text{FCE}}$ thus reflects the shared visual characteristics across all single-instance proposals. By grounding exemplar selection in feature-level consensus rather than count statistics, FCE complements DGE by emphasizing visual coherence.
Ultimately, $\mathbf{b}_{\text{DAE}}$, $\mathbf{b}_{\text{DGE}}$, and $\mathbf{b}_{\text{FCE}}$ constitute the final exemplar set, which is jointly provided as multiple support exemplars to the density estimator to produce the final density map and count prediction. Selecting one exemplar per stage, rather than all exemplars from a single stage, introduces cross-perspective diversity (see ablations in Sec. 4).
@@ 4 Experiments
Datasets and Evaluation Metrics. We evaluate our CountZES on diverse benchmarks including FSC-147 (Shi et al., 2022), CARPK (Hsieh et al., 2017), PerSense-D (Siddiqui et al., 2025), and two medical cell-counting datasets, VGG (Lempitsky and Zisserman, 2010) and MBM (Paul Cohen et al., 2017). See Appendix. A.4 for dataset details. Following prior ZOC works, we employ mean absolute error (MAE) to measure counting accuracy and root mean square error (RMSE) to assess model robustness.
@@
Implementation Details. CountZES employs CLIP (Radford et al., 2021) for text–image alignment and GroundingDINO (Liu et al., 2024) for initial box proposals, with the detection threshold set to 0.15. SAM (Kirillov et al., 2023) is used for mask generation leveraging point prompts across the SSES, GGES, and FRES modules. To demonstrate the model-agnostic nature of CountZES, we utilize CounTR (Liu et al., 2022) and DSALVANet (He et al., 2024) as pretrained off-the-shelf density estimators (DE). We distinguish between in-domain and cross-domain settings: in-domain uses a DE pretrained on the same dataset as evaluation, while cross-domain employs DE pretrained on a dataset different from the one being tested.
Quantitative Results on FSC-147. We evaluate CountZES under two configurations, as summarized in Table 1. In the in-domain setting, DSALVANet pretrained on FSC-147 is employed as the DE. Under this configuration, CountZES achieves the lowest RMSE and a competitive MAE among ZOC-optimized methods, despite being an inference-only framework. Compared to the leading exemplar selection–based approach, VA-Count (Zhu et al., 2024), CountZES reduces MAE by 11.8% and RMSE by 29.3%, demonstrating the effectiveness of our exemplar selection strategy. For comparison with inference-only approaches, we further evaluate CountZES in a cross-domain setting using DE pretrained on CARPK. Among inference-only methods, CountZES attains the lowest MAE and RMSE.
@@
Table 1: Comparison of CountZES with SOTA ZOC methods on FSC-147 dataset. Best results are in bold and second-best are underlined. indicates ZOC methods based on exemplar discovery. * indicates that CountingDINO is not zero-shot (box prompts).
Method Venue Inference MAE RMSE
only
ZSOC (Xu et al., 2023) CVPR’23 22.09 115.17
CLIP-Count (Jiang et al., 2023) ACM’23 17.78 106.62
CounTX (Amini et al., 2023) BMVC’23 15.88 106.29
VLCounter (Kang et al., 2024) AAAI’24 17.05 106.16
PseCo (Huang et al., 2024) CVPR’24 16.58 129.77
DAVE (Pelhan et al., 2024b) ICCV’24 14.90 103.42
VA-Count (Zhu et al., 2024) ECCV’24 17.88 129.31
CountGD (Amini et al., 2024) NeurIPS’24 14.76 120.42
GeCo (Pelhan et al., 2024a) NeurIPS’24 13.30 108.72
T2ICount (Qian et al., 2025) CVPR’25 11.76 97.86
CountZES (in-domain) (Ours) 15.77 91.40
CountingDINO* (Pacini et al., 2026) WACV’26 20.93 71.37
SAM (Kirillov et al., 2023) ICCV’23 42.48 137.50
Count Anything (Ma et al., 2023) arXiv’23 27.97 131.24
G-DINO (Liu et al., 2024) ECCV’24 59.23 159.28
TFOC (Shi et al., 2024) WACV’24 24.79 137.15
TFCAC (Lin et al., 2025) WACV’25 23.59 113.60
OmniCount (Mondal et al., 2025) AAAI’25 21.46 133.28
CountZES (Ours) 21.09 110.14
@@
Method Inference-only MAE RMSE
CLIP-Count (Jiang et al., 2023) 11.96 16.61
CounTX (Amini et al., 2023) 8.13 10.87
G-DINO (Liu et al., 2024) 29.72 31.60
VLCounter (Kang et al., 2024) 6.46 8.68
VA-Count (Zhu et al., 2024) 10.63 13.20
GeCo (Pelhan et al., 2024a) 10.34 14.73
CountGD (Amini et al., 2024) 3.83 5.41
T2ICount (Qian et al., 2025) 8.61 13.47
TFOC (Shi et al., 2024) 14.35 17.22
OmniCount (Mondal et al., 2025) 13.41 16.85
CountingDINO (Pacini et al., 2026) 21.26 28.20
CountZES (Ours) 7.24 9.22
Quantitative Results on CARPK. We evaluate CountZES on CARPK under a cross-domain setting using DSALVANet pretrained on FSC-147 as the density estimator. As shown in Table 2, CountZES achieves the lowest MAE and RMSE among inference-only approaches, while remaining competitive with ZOC-optimized methods that rely on task-specific training. We attribute this strong generalization to the proposed multi-stage exemplar selection strategy, which selects representative single-instance exemplars by leveraging complementary semantic, statistical, and feature-level cues. This principled exemplar selection without task-specific training enables robust counting even under substantial domain shift.
@@
Density-based Class-wise
Method (Inf-only) Low Med High Overall
MAE RMSE MAE RMSE MAE RMSE MAE RMSE
GeCo (✗) 11.63 17.84 10.30 20.61 12.38 24.13 13.12 28.36
T2ICount (✗) 30.04 122.91 25.16 76.95 20.62 52.23 24.95 88.21
TFOC (✓) 7.20 10.36 16.07 21.63 36.88 63.82 17.79 35.74
CountZES (✓) 6.86 19.22 10.26 14.44 20.36 43.65 12.29 28.32
@@
Method (Inf-only) VGG MBM
MAE RMSE MAE RMSE
GeCo (✗) 67.32 77.53 92.59 104.52
T2ICount (✗) 151.35 162.81 30.01 44.44
TFOC (✓) 144.91 158.52 64.25 75.67
CountZES (✓) 45.55 74.41 22.16 40.06
Figure 6: Qualitative comparison of CountZES with T2ICount (Qian et al., 2025) and GeCo (Pelhan et al., 2024a) across diverse scenarios, including natural (FSC-147 and PerSense-D), aerial (CARPK), and medical cell-counting (VGG and MBM) images.
Quantitative Results on PerSense-D. Following the same cross-dataset evaluation protocol, we evaluate CountZES on the PerSense-D dataset using CounTR pretrained on FSC147 (Table 3). In addition to overall category-wise performance, we conduct fine-grained analysis across three density-based splits; low, medium, and high, as defined in (Siddiqui et al., 2025). CountZES consistently achieves competitive performance across varying crowding conditions, outperforming the inference-only approach TFOC and remaining comparable to ZOC-optimized methods. Notably, 16 out of 28 classes ($\approx$57%) in PerSense-D overlap with FSC-147, but still T2ICount, which achieves strong in-domain performance on FSC-147, struggles on PerSense-D. We attribute this degradation to the reliance on textual supervision during training in T2ICount, which couples linguistic priors with numerical estimation. Such entanglement is particularly brittle under distribution shift, where object appearance, density, and scale differ substantially across datasets. In contrast, CountZES employs textual input solely as an inference-time semantic probe to guide exemplar discovery, while the counting backbone remains purely visual and exemplar-driven. By decoupling textual guidance from numerical estimation, it mitigates text-induced bias and preserves semantic flexibility, thereby improving robustness to distributional shifts for the same category across datasets.
Quantitative Results on VGG and MBM. To further assess the cross-domain generalization capability of CountZES, we evaluate it on two medical-domain cell counting datasets: VGG and MBM. As summarized in Table 4, CountZES achieves remarkable cross-domain performance, outperforming both inference-only and ZOC-optimized approaches by a substantial margin.
Qualitative Result & Runtime Efficiency. Fig. 6 presents a qualitative comparison between CountZES and SOTA methods on diverse images, spanning natural, aerial, and medical domains. On average, CountZES infers the total count in 2.3 seconds per image and achieves the best performance–runtime tradeoff among inference-only approaches (see Figure. 11 and Appendix A.6).
Ablation Study. Since each stage builds upon the exemplar generated by the previous one, we perform a progressive ablation to quantify the incremental contribution of each stage toward overall counting performance. As shown in Table 5, adding successive stages consistently improves accuracy, validating their complementary roles. Finally, to examine the diversity induced by each stage in exemplar selection, we conduct another ablation under a consistent 3-exemplar setting by selecting the top three exemplars from each stage and comparing them against the standard CountZES configuration that uses one exemplar per stage (Table 5). The latter achieves superior performance (Figure. 7), indicating that cross-stage collaboration yields a richer, complementary, and more balanced exemplar set than any stage in isolation. We discuss the limitations of CountZES in Appendix. A.7.
@@
CountZES 15.77 91.40